# ClutterDexGrasp: A Sim-to-Real System for General Dexterous Grasping in Cluttered Scenes

### v2

@@ B.4 Safety Curriculum
3:     $\tau=\text{rollout\_policy}(\pi_{\theta})$triangleright get rollout trajectory
4:     $\pi_{\theta}=\text{optimize\_policy}(\pi_{\theta},\tau)$triangleright update policy
@@
7:         $w=\text{evaluate\_policy}(\pi_{\theta})$triangleright get success rate $w$
@@
10:              $f=\min(f+\Delta f,f_{\max})$triangleright tighten safety constraint
@@ B.6 Training Details
All the training and experiment in the paper were run on a single GeForce RTX 4090 GPU with i9-13900K CPU. The hyperparameters for teacher policy training are listed in Tab. 6.
@@ C.3 Point Cloud Observation Processing
The point cloud is created from a single-view depth image. The point-cloud pre-processing includes four steps: (i) cropping the point cloud to the workspace region using a manually defined bounding box, (ii) downsampling the point cloud to 3584 points, (iii) transforming the point cloud from the camera frame to the robot base frame, and (iv) replacing the table surface with a synthetic point cloud (512 points) to mitigate point cloud holes caused by the flat table. However, we found that adding Gaussian noise and random transformation to point cloud during DP3 training did not improve policy performance in real-world settings.
@@ C.4 System Identification
To minimize the sim-to-real gap, we calibrate the physical and dynamic parameters of robot via system identification (SI) by aligning simulated trajectories with their real-world counterparts, following the approach in [42].
@@ C.5 Training Details
The hyperparameters for student policy training are listed in Tab. 6.
Appendix D Simulation Environment Details
@@ E.1 Real-World Hardware Setup
To improve deployment stability, we introduce several practical adjustments. Since the system occasionally failed to detect successful grasps, leading to redundant actions, we use finger torque as a success signal: Once it exceeds a threshold, the grasp is considered successful and a predefined lift-up action is executed. For safety, we add a threshold on end-effector force; if exceeded, the end-effector is gently raised avoiding and avoid approaching dangerous force boundaries. Additionally, we slow down the end-effector slightly compared to simulation to ensure smoother and safer real-world execution.

### v3

@@ B.4 Safety Curriculum
3:  $\tau=\text{rollout\_policy}(\pi_{\theta})$ $\triangleright$ get rollout trajectory
4:  $\pi_{\theta}=\text{optimize\_policy}(\pi_{\theta},\tau)$ $\triangleright$ update policy
@@
7:   $w=\text{evaluate\_policy}(\pi_{\theta})$ $\triangleright$ get success rate $w$
@@
10:     $f=\min(f+\Delta f,f_{\max})$ $\triangleright$ tighten safety constraint
@@ B.6 Training Details
All the training and experiment in the paper were run on a single GeForce RTX 4090 GPU with i9-13900K CPU. Memory usage and training time are summarized in Table 7, and PPO hyperparameters for the teacher policy are listed in Table 7.
@@
Mem. Time
Teacher 22G 8 days
Student 22G 1 day
Table 6: Computation resources.
@@ C.3 Point Cloud Observation Processing
The point cloud is created from a single-view depth image. The point-cloud pre-processing includes four steps: (i) cropping the point cloud to the workspace region using a manually defined bounding box, (ii) downsampling the point cloud to 3584 points, (iii) transforming the point cloud from the camera frame to the robot base frame, and (iv) replacing the table surface with a synthetic point cloud (512 points) to mitigate point cloud holes caused by the flat table. We apply consistent point cloud preprocessing across simulation and real-world data to ensure alignment. However, we found that adding Gaussian noise and random transformation to point cloud during DP3 training did not improve policy performance in real-world settings.
@@ C.4 System Identification
To minimize the sim-to-real gap, we calibrate the physical and dynamic parameters of robot via system identification (SI) by aligning simulated trajectories with their real-world counterparts, following the approach in [60, 42].
@@ C.5 Training Details
Memory usage and training time are summarized in Table 6, and the hyperparameters for student policy training are listed in Tab. 7.
Appendix D Simulation Environment Generation Details
@@ E.1 Real-World Hardware Setup
To improve deployment stability, we introduce several practical adjustments. Since the system occasionally failed to detect successful grasps, leading to redundant actions, we use finger torque as a success signal: Once it exceeds a threshold, the grasp is considered successful and a predefined lift-up action is executed. To guarantee safety, we add a threshold on end-effector force; if exceeded, the end-effector is gently raised to avoid approaching dangerous force boundaries. Additionally, we slow down the end-effector slightly compared to simulation to ensure smoother and safer real-world execution. Note that, in all real-world experiments, no safety-related human intervention was observed.